# A note on restricted partition functions of Pushpa and Vasuki

### v1

@@ Abstract.
We establish infinite families of congruences modulo arbitrary powers of 22 for three restricted partition functions $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ recently introduced by Pushpa and Vasuki by employing elementary $q$-series techniques.
@@ 1. Introduction
Pushpa and Vasuki [9] introduced in 2022 seven restricted partition functions, three of which have the following generating functions defined by
@@
and proved some congruences modulo 22 and 55 for $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$. Consequently, Dasappa et. al. [4] obtained more congruences and linear identities for these restricted partition functions. Very recently, Nath and Saikia [8] established infinite families of congruences modulo powers of 55 for $M(n)$ and $T^{\ast}(n)$ and particular congruences for $P^{\ast}(n)$, and settled the conjecture of Dasappa et. al. [4, Conjecture 1] by relying on the Mathematica package RaduRK created by Smoot [12], which is based on Radu’s Ramanujan–Kolberg algorithm [10].
In this short note, we explore more arithmetic properties of $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ using elementary $q$-series techniques. Our main results reveal infinite families of congruences modulo powers of 22 for these restricted partition functions, which generalize some results of Nath and Saikia [8].

### v2

# Congruences modulo powers of 22 for three restricted partition functions of Pushpa and Vasuki

@@ Abstract.
We establish infinite families of congruences modulo arbitrary powers of 22 for the three restricted partition functions $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ introduced by Pushpa and Vasuki by employing elementary $q$-series techniques. These generalize some particular congruences for $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ recently found by Nath and Saikia.
@@ 1. Introduction
Pushpa and Vasuki [9] introduced in 2022 three restricted partition functions $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ defined as follows:
  1. (1)
    $M(n)$ is the difference $m_{(d,e)}(n)-m_{(d,o)}(n)$, where $m_{(d,e)}(n)$ (respectively, $m_{(d,o)}(n)$) is the number of partitions of $n$ such that
    • parts that are congruent to $\pm 1,\pm 3\pmod{10}$ have one color,
    • parts that are congruent to $0\pmod{10}$ are distinct and have eight colors,
    • parts that are not congruent to $0,\pm 1,\pm 3\pmod{10}$ are distinct and have four colors each, and
    • the number of parts that are not congruent to $\pm 1,\pm 3\pmod{10}$ are even (respectively, odd).
  2. (2)
    $T^{\ast}(n)$ is the difference $t_{(d,e)}(n)-t_{(d,o)}(n)$, where $t_{(d,e)}(n)$ (respectively, $t_{(d,o)}(n)$) is the number of distinct partitions of $n$ with even (respectively, odd) number of parts such that
    • parts that are congruent to $0\pmod{10}$ have eight colors,
    • parts that are congruent to $\pm 1,\pm 3\pmod{10}$ have five colors each, and
    • parts that are congruent to $\pm 2,\pm 4,5\pmod{10}$ have four colors each.
  3. (3)
    $P^{\ast}(n)$ is the difference $p_{(d,e)}(n)-p_{(d,o)}(n)$, where $p_{(d,e)}(n)$ (respectively, $p_{(d,o)}(n)$) is the number of distinct partitions of $n$ with even (respectively, odd) number of parts such that
    • parts that are congruent to $0\pmod{5}$ have eight colors, and
    • parts that are not congruent to $0\pmod{5}$ have four colors each.
Pushpa and Vasuki [9] obtained the generating functions for $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ given by
@@
and proved some congruences modulo 22 and 55 for $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$. Subsequently, Dasappa et. al. [4] found more congruences and linear identities for these restricted partition functions. Very recently, Nath and Saikia [8] established infinite families of congruences modulo powers of 55 for $M(n)$ and $T^{\ast}(n)$ and particular congruences for $P^{\ast}(n)$. They also settled the conjecture of Dasappa et. al. [4, Conjecture 1] by relying on the Mathematica package RaduRK created by Smoot [12], which is based on Radu’s Ramanujan–Kolberg algorithm [10].
We explore in this note more arithmetic properties of $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ using elementary $q$-series techniques. Our main results reveal infinite families of congruences modulo powers of 22 for these restricted partition functions, which generalizes some of the results of Nath and Saikia [8].